# Seg2Any: Open-set Segmentation-Mask-to-Image Generation with Precise Shape and Semantic Control

### v1

@@ 1 Introduction
Layout-to-image generation has been proposed and designed to generate images based on specified layout conditions, including spatial locations and descriptions of entities. These layout conditions come in various forms, such as bounding boxes [19, 50, 36, 45, 22], depth maps [52, 53, 51], segmentation masks [39, 21, 44], etc. This paper focuses on segmentation-mask-to-image (S2I) generation, where segmentation masks dictate the spatial locations of entities, and text descriptions specify their semantic content, thereby enabling the most fine-grained control over the images.
@@
Existing S2I methods can be mainly divided into two categories: \@slowromancapi@) Methods that integrate segmentation masks as additional conditional inputs, such as ControlNet [46], ControlNet++ [18] and T2I-Adapter [23]. These methods often fail to align regional textual descriptions with their respective regions, resulting in semantic inconsistency in the generated images (see Figure  2 (a)). \@slowromancapii@) Methods based on the masked attention mechanism, which restricts each text embedding to attend solely to the respective image embeddings (e.g. FreestyleNet [39], PLACE [21] and EliGen [44]). Although these methods achieve semantic alignment, they fall short in precise shape preservation, as shown in Figure 2 (b) and (c). We attribute this shape inconsistency to the loss of spatial information when segmentation masks are compressed into the latent space. Notably, compared to the UNet [31] architecture with 8×8$\times$8 × downsampling, recent advanced DiT [25] architectures employ a more aggressive 16×16$\times$16 × downsampling, which further amplifies the loss of spatial information.
@@
To address these challenges, we propose Seg2Any, a novel S2I framework built upon advanced multimodal diffusion transformers (e.g. FLUX [17]). Seg2Any mainly relies on two key innovations: \@slowromancapi@) Semantic-Shape Decoupled Layout Conditions Injection. We decouple segmentation mask conditions into two components: the shape condition (high-frequency shape information) and the regional semantic condition (low-frequency semantic information). For the injection of regional semantic conditions, we employ a Semantic Alignment Attention Mask. This mechanism tightly binds each entity’s image tokens to its corresponding text prompts, ensuring that the generated image is semantically consistent with the input descriptions at the entity level. For the injection of shape condition, we propose Sparse Shape Feature Adaptation to integrate key spatial structure into the model efficiently. Conventional approaches [46, 18] assign fixed colors to distinct categories in the semantic segmentation maps, which is only suitable for closed-set S2I generation and cannot be generalized to open-set scenarios. In contrast, we introduce an Entity Contour Map as our category-agnostic shape representation, which consists of entity contours extracted from segmentation masks. This representation is inherently sparse, preserving essential shape details in a compact and efficient manner. Following OminiControl [34], we integrate the condition tokens with text and noisy image tokens into a unified sequence, allowing them to interact directly through multi-modal attention [9]. As shown in Figure 2 (d), our approach achieves both semantic and shape consistency simultaneously.
@@
\@slowromancap
ii@) Attribute Isolation via Image Self-Attention Mask. Multi-instance generation often suffers from attribute leakage, where the visual attributes of one entity may affect others. To mitigate this problem, we introduce an Attribute Isolation Attention Mask strategy that prevents cross-entity information leakage by ensuring that image tokens associated with each entity are isolated from those of others.
Current dense caption datasets [30, 42, 28, 8] are often limited by their closed-set vocabularies and coarse-grained descriptions, which constrain their effectiveness in training S2I models. Recent advances in open-source vision language models (VLMs), such as Qwen2-VL-72B [35], have significantly reduced the performance gap with close-source VLMs like GPT-4V [1], making it feasible to create large-scale and richly annotated datasets. Leveraging the capabilities of Qwen2-VL-72B, we construct Segment Anything with Captions 1 Million (SACap-1M), a large-scale dataset derived from the diverse and high-resolution SA-1B dataset [15]. SACap-1M contains 1 million image-text pairs and 5.9 million segmented entities, each comprised of a segmentation mask and a detailed regional caption, with captions averaging 58.6 words per image and 14.1 words per entity. We further present the SACap-Eval, a benchmark for assessing the quality of open-set S2I generation.
@@ 2.1 Text-to-Image Generation
Text-to-image (T2I) generation [31, 4, 9] has undergone significant advancement in recent years. Motivated by advances in large-scale transformer architectures, the Diffusion Transformer (DiT) [25] was introduced. Building upon this foundation, recent models like SD3 [9] and FLUX [17] further propose the Multimodal Diffusion Transformer (MM-DiT), which treats text as an independent modality and incorporates flow matching objectives, achieving state-of-the-art results. To integrate additional condition images (e.g., canny maps, depth maps, subject references) into MM-DiT, OminiControl [34] introduces a novel controllable framework. It concatenates condition image, text, and noisy image tokens, and then employs task-specific LoRA [13] modules to handle various conditions within a unified pipeline while maintaining minimal trainable parameters.
@@ 2.2 Layout-to-Image Generation
Layout-to-Image (L2I) generation is a task that synthesizes images guided by spatial layout conditions and entity-level textual descriptions. Existing L2I methods can be categorized according to the type of layout condition they employ. These include: Bounding box-based approaches [19, 50, 36, 45, 22, 11, 10, 33, 7, 48, 40] typically use rectangular spatial constraints to guide the image generation process, offering a coarse-grained layout control. Depth map-based approaches [52, 53, 51] utilize depth maps to achieve fine-grained spatial control similar to segmentation masks. For instance, 3DIS [52] divides multi-instance generation into two stages: it first generates a depth map via a text-to-depth model, then uses a pre-trained depth-to-image model to synthesize images with multi-instance attribute control. DreamRenderer [51], a training-free approach based on pre-trained controllable T2I models (FLUX.1-Depth [17] and FLUX.1-Canny [17]), identifies that middle layers in the FLUX model are responsible for instance-level rendering while shallow and deep layers capture global context. Consequently, it applies a hard image self-attention mask only to the middle layers to prevent attribute leakage. However, as a training-free approach, directly applying the attention mask severely impairs overall visual harmony.
Segmentation mask-based approaches [39, 21, 44, 24, 14, 2, 43] focus on pixel-level layout control. For example, FreestyleNet [39] uses binary attention weights in the cross-attention module that assigns a value $1$ to allow text tokens to bind to corresponding image regions and 00 to prevent attention from unrelated areas. However, this approach requires downsampling the segmentation masks to align with the lower-resolution latent features, thereby sacrificing spatial detail. To alleviate the above issue, PLACE [21] introduces a layout control map that softens the attention weights. Instead of hard binary assignments in FreestyleNet, it calculates the area proportions of different entities within the receptive field of every image token, yielding soft attention weights. Yet, this approach struggles to mitigate spatial detail loss in MM-DiT architectures, where 16× downsampling (compared to 8× in U-Net) leads to more severe degradation of spatial information. The work most similar to ours is EliGen [44], which is also built on FLUX. Unlike our approach, it is trained on datasets with bbox annotations and only supports loose position control through scribble-style masks, whereas our approach enables both strict and loose mask position control. Additionally, it relies solely on a masked attention mechanism without injecting explicit spatial information.
@@ 3.1 Problem Definition
Figure 3: (a) An overview of the Seg2Any framework. Segmentation masks are transformed into Entity Contour Map, then encoded as condition tokens via frozen VAE. Negligible tokens are filtered out for efficiency. The resulting text, image, and condition tokens are concatenated into a unified sequence for MM-Attention. Our framework applies LoRA to all branches, achieving S2I generation with minimal extra parameters. (b) Attention Masks in MM-Attention, including Semantic Alignment Attention Mask (Section 3.2.1) and Attribute Isolation Attention Mask (Section 3.3).
@@ 3.2.2 Sparse Shape Feature Adaptation
The factor $\gamma\in(0,1]$ serves to modulate the strength of shape conditioning. As $\gamma$ approaches $1$, $\log(\gamma)$ tends to 00, and the condition tokens retain full influence, enforcing strict adherence to the shape guidance. In contrast, as $\gamma$ approaches 00, $\log(\gamma)$ tends to $\infty$- ∞, which suppresses the contribution of condition tokens, thus making scribble-style segment masks feasible for flexible control.
@@ 3.4 SACap-1M Dataset
To address the lack of large-scale and fine-grained datasets for S2I generation, we construct SACap-1M, containing 1 million image-text pairs and 5.9 million segmented entities with detailed descriptions. We propose an automated pipeline for data annotation and filtering: \@slowromancapi@) Image Filtering. We select images from the high-resolution and wide-ranging SA-1B [15] dataset. Initially, we remove images that are excessively large or small in size, as well as those with extreme aspect ratios. Subsequently, we apply the LAION-Aesthetics predictor [32] to filter out low-quality images with an aesthetic score below $5$@slowromancapii@) Entity Extraction. The SA-1B dataset provides accurate, category-agnostic masks for each image. However, each image contains on average over 100 masks, many of which are nested. To filter masks, we retain only top-level masks by removing those that are contained within top-level masks. Additionally, we discard masks whose area is smaller than $1\%$ of the total image area. Finally, we exclude images whose number of remaining masks falls outside the range of 1 to 20. \@slowromancapiii@) Regional and Global Caption Annotation. We employ the open-source Qwen2-VL-72B [35] model to generate captions for each entity and the global image, yielding an average of 58.6 words per image and 14.1 words per entity. See supplemental materials for more details.
@@ 4.2 Quantitative Results
Method Class-agnostic MIoU $\uparrow$ Region-wise Quality Global-wise Quality
@@ 4.3 Ablation Study
We conduct an ablation study to evaluate the individual contributions of our key components on the SACap-Eval benchmark (see Table 4). Introducing explicit shape condition via Sparse Shape Feature Adaptation (SSFA) significantly improves the class-agnostic MIoU metric over using only the Semantic Alignment Attention Mask (SAA), confirming the effectiveness of preserving shape details. Further, adding the Attribute Isolation Attention Mask (AIA) notably improves region-wise quality, indicating better prevention of attribute leakage. The training-free version degrades global quality; however, after training on our large-scale dataset, the overall image quality improves significantly. Table 4 also shows that incorporating Condition Image Token Filtering (CITF) leaves performance unchanged; its impact on computational cost is further analyzed in the supplemental materials.
On COCO-Stuff and ADE20K, we observe a similar trend: SSFA substantially improves MIoU, and CITF does not affect performance (see Table 3).
@@
[…]
@@ Appendix A More Details of Condition Image Token Filtering
We analyze the computational cost of Condition Image Token Filtering (CITF) on three benchmark datasets: COCO-Stuff, ADE20K, and SACap-1M. For each dataset, we consider five settings to comprehensively analyze the computational cost: \@slowromancapi@) without condition tokens, representing the lowest computation. \@slowromancapii@–\@slowromancapiv@) CITF applied with the minimum, maximum, and average numbers of retained condition tokens, respectively. \@slowromancapv@) without CITF, i.e., no token filtering, representing the highest computation. For each configuration, we report the resulting average image generation time and multiply–accumulate operations (MACs).
@@ Appendix B More Details of Shape Guidance Strength Modulation
Appendix C More Details on Datasets and Benchmarks
C.1 Comparison with Existing Dense Caption Datasets
Table 5 presents a detailed comparison between our SACap-1M dataset and several recent dense caption datasets. Unlike previous datasets such as PixelLM-MUSE [30], Osprey [42], and COCONut-PanCap [8] which are constructed upon closed-set label spaces (e.g., LVIS [12] and COCO [20]), our proposed SACap-1M dataset provides open-set segmented entities, enabling much broader generalization and flexibility in open-set segmentation-mask-to-image generation. Compared to GLaMM-GranD [28], also derived from SA-1B [15], our SACap-1M achieves significantly higher segmentation mask density and caption granularity by employing the advanced open-source Qwen2-VL-72B [35] vision-language model to generate more accurate and fine-grained regional captions.

### v3

@@ 1 Introduction
Layout-to-image generation has been proposed and designed to generate images based on specified layout conditions, including spatial locations and descriptions of entities. These layout conditions come in various forms, such as bounding boxes [22, 60, 44, 19, 54, 25], depth maps [62, 63, 61], segmentation masks [47, 24, 53], etc. This paper focuses on segmentation-mask-to-image (S2I) generation, where segmentation masks dictate the spatial locations of entities, and text descriptions specify their semantic content, thereby enabling the most fine-grained control over the images.
@@
Existing S2I methods can be mainly divided into two categories: I) Methods that integrate segmentation masks as additional conditional inputs, such as ControlNet [56], ControlNet++ [20] and T2I-Adapter [26]. These methods often fail to align regional textual descriptions with their respective regions, resulting in semantic inconsistency in the generated images (see Figure  2 (a)). II) Methods based on the masked attention mechanism, which restricts each text embedding to attend solely to the respective image embeddings (e.g. FreestyleNet [47], PLACE [24] and EliGen [53]). Although these methods achieve semantic alignment, they fall short in precise shape preservation, as shown in Figure 2 (b) and (c). We attribute this shape inconsistency to the loss of spatial information when segmentation masks are compressed into the latent space. Notably, compared to the UNet [34] architecture with $8\times$ downsampling, recent advanced DiT [28] architectures employ a more aggressive $16\times$ downsampling, which further amplifies the loss of spatial information.
@@
To address these challenges, we propose Seg2Any, a novel S2I framework built upon advanced multimodal diffusion transformers (e.g. FLUX [18]). Seg2Any mainly relies on two key innovations: I) Semantic-Shape Decoupled Layout Conditions Injection. We decouple segmentation mask conditions into two components: the shape condition (high-frequency shape information) and the regional semantic condition (low-frequency semantic information). For the injection of regional semantic conditions, we employ a Semantic Alignment Attention Mask. This mechanism tightly binds each entity’s image tokens to its corresponding text prompts, ensuring that the generated image is semantically consistent with the input descriptions at the entity level. For the injection of shape condition, we propose Sparse Shape Feature Adaptation to integrate key spatial structure into the model efficiently. Conventional approaches [56, 20] assign fixed colors to distinct categories in the semantic segmentation maps, which is only suitable for closed-set S2I generation and cannot be generalized to open-set scenarios. In contrast, we introduce an Entity Contour Map as our category-agnostic shape representation, which consists of entity contours extracted from segmentation masks. This representation is inherently sparse, preserving essential shape details in a compact and efficient manner. Following OminiControl [37], we integrate the condition tokens with text and noisy image tokens into a unified sequence, allowing them to interact directly through multi-modal attention [9]. As shown in Figure 2 (d), our approach achieves both semantic and shape consistency simultaneously.
@@
II) Attribute Isolation via Image Self-Attention Mask. Multi-instance generation often suffers from attribute leakage, where the visual attributes of one entity may affect others. To mitigate this problem, we introduce an Attribute Isolation Attention Mask strategy that prevents cross-entity information leakage by ensuring that image tokens associated with each entity are isolated from those of others.
Current dense caption datasets [33, 51, 31, 8, 50] are often limited by their closed-set vocabularies and coarse-grained descriptions, which constrain their effectiveness in training S2I models. Recent advances in open-source vision language models (VLMs), such as Qwen2-VL-72B [43], have significantly reduced the performance gap with close-source VLMs like GPT-4V [1], making it feasible to create large-scale and richly annotated datasets. Leveraging the capabilities of Qwen2-VL-72B, we construct Segment Anything with Captions 1 Million (SACap-1M), a large-scale dataset derived from the diverse and high-resolution SA-1B dataset [16]. SACap-1M contains 1 million image-text pairs and 5.9 million segmented entities, each comprised of a segmentation mask and a detailed regional caption, with captions averaging 58.6 words per image and 14.1 words per entity. We further present the SACap-Eval, a benchmark for assessing the quality of open-set S2I generation.
@@ 2.1 Text-to-Image Generation
Text-to-image (T2I) generation [34, 4, 9, 42, 38] has undergone significant advancement in recent years. Motivated by advances in large-scale transformer architectures, the Diffusion Transformer (DiT) [28] was introduced. Building upon this foundation, recent models like SD3 [9] and FLUX [18] further propose the Multimodal Diffusion Transformer (MM-DiT), which treats text as an independent modality and incorporates flow matching objectives, achieving state-of-the-art results. To integrate additional condition images (e.g., canny maps, depth maps, subject references) into MM-DiT, OminiControl [37] introduces a novel controllable framework. It concatenates condition image, text, and noisy image tokens, and then employs task-specific LoRA [13] modules to handle various conditions within a unified pipeline while maintaining minimal trainable parameters.
@@ 2.2 Layout-to-Image Generation
Layout-to-Image (L2I) generation is a task that synthesizes images guided by spatial layout conditions and entity-level textual descriptions. Existing L2I methods can be categorized according to the type of layout condition they employ. These include: Bounding box-based approaches [22, 60, 44, 19, 54, 55, 25, 11, 10, 36, 7, 58, 48] typically use rectangular spatial constraints to guide the image generation process, offering a coarse-grained layout control. Depth map-based approaches [62, 63, 61] utilize depth maps to achieve fine-grained spatial control similar to segmentation masks. For instance, 3DIS [62] divides multi-instance generation into two stages: it first generates a depth map via a text-to-depth model, then uses a pre-trained depth-to-image model to synthesize images with multi-instance attribute control. DreamRenderer [61], a training-free approach based on pre-trained controllable T2I models (FLUX.1-Depth [18] and FLUX.1-Canny [18]), identifies that middle layers in the FLUX model are responsible for instance-level rendering while shallow and deep layers capture global context. Consequently, it applies a hard image self-attention mask only to the middle layers to prevent attribute leakage. However, as a training-free approach, directly applying the attention mask severely impairs overall visual harmony.
Segmentation mask-based approaches [47, 24, 53, 27, 15, 2, 52] focus on pixel-level layout control. For example, FreestyleNet [47] uses binary attention weights in the cross-attention module that assigns a value $1$ to allow text tokens to bind to corresponding image regions and 0 to prevent attention from unrelated areas. However, this approach requires downsampling the segmentation masks to align with the lower-resolution latent features, thereby sacrificing spatial detail. To alleviate the above issue, PLACE [24] introduces a layout control map that softens the attention weights. Instead of hard binary assignments in FreestyleNet, it calculates the area proportions of different entities within the receptive field of every image token, yielding soft attention weights. Yet, this approach struggles to mitigate spatial detail loss in MM-DiT architectures, where 16× downsampling (compared to 8× in U-Net) leads to more severe degradation of spatial information. The work most similar to ours is EliGen [53], which is also built on FLUX. Unlike our approach, it is trained on datasets with bbox annotations and only supports loose position control through scribble-style masks, whereas our approach enables both strict and loose mask position control. Additionally, it relies solely on a masked attention mechanism without injecting explicit spatial information.
@@ 3.1 Problem Definition
Figure 3: (a) An overview of the Seg2Any framework. Segmentation masks are transformed into Entity Contour Map, then encoded as condition tokens via frozen VAE. Negligible tokens are filtered out for efficiency. The resulting text, image, and condition tokens are concatenated into a unified sequence for $M$-Attention. Our framework applies LoRA to all branches, achieving S2I generation with minimal extra parameters. (b) Attention Masks in $M$-Attention, including Semantic Alignment Attention Mask (Section 3.2.1) and Attribute Isolation Attention Mask (Section 3.3).
@@ 3.2.2 Sparse Shape Feature Adaptation
The factor $\gamma\in(0,1]$ serves to modulate the strength of shape conditioning. As $\gamma$ approaches $1$, $\log(\gamma)$ tends to 0, and the condition tokens retain full influence, enforcing strict adherence to the shape guidance. In contrast, as $\gamma$ approaches 0, $\log(\gamma)$ tends to -$\infty$, which suppresses the contribution of condition tokens, thus making scribble-style segment masks feasible for flexible control.
@@ 3.4 SACap-1M Dataset
To address the lack of large-scale and fine-grained datasets for S2I generation, we construct SACap-1M, containing 1 million image-text pairs and 5.9 million segmented entities with detailed descriptions. We propose an automated pipeline for data annotation and filtering: I) Image Filtering. We select images from the high-resolution and wide-ranging SA-1B [16] dataset. Initially, we remove images that are excessively large or small in size, as well as those with extreme aspect ratios. Subsequently, we apply the LAION-Aesthetics predictor [35] to filter out low-quality images with an aesthetic score below $5$. II) Entity Extraction. The SA-1B dataset provides accurate, category-agnostic masks for each image. However, each image contains on average over 100 masks, many of which are nested. To filter masks, we retain only top-level masks by removing those that are contained within top-level masks. Additionally, we discard masks whose area is smaller than $1\%$ of the total image area. Finally, we exclude images whose number of remaining masks falls outside the range of 1 to 20. III) Regional and Global Caption Annotation. We employ the open-source Qwen2-VL-72B [43] model to generate captions for each entity and the global image, yielding an average of 58.6 words per image and 14.1 words per entity. See supplemental materials for more details.
@@ 4.2 Quantitative Results
Method
Class-agnostic
MIoU $\uparrow$
Region-wise Quality Global-wise Quality
@@ 4.3 Ablation Study
We conduct an ablation study to evaluate the individual contributions of our key components on the SACap-Eval benchmark (see Table 4). Introducing explicit shape condition via Sparse Shape Feature Adaptation (SSFA) significantly improves the class-agnostic MIoU metric over using only the Semantic Alignment Attention Mask (SAA), confirming the effectiveness of preserving shape details. Further, adding the Attribute Isolation Attention Mask (AIA) notably improves region-wise quality, indicating better prevention of attribute leakage. The training-free version degrades global quality; however, after training on our large-scale dataset, the overall image quality improves significantly. Table 4 also shows that incorporating Condition Image Token Filtering (CITF) leaves performance almost unchanged; its impact on computational cost is further analyzed in the supplemental materials.
On COCO-Stuff and ADE20K, we observe a similar trend: SSFA substantially improves MIoU, and CITF shows a marginal impact on performance. (see Table 3).
@@
SAA SSFA $\dagger AIA\ddagger$ CITF Spatial $\uparrow$ Color $\uparrow$ Shape $\uparrow$ Texture $\uparrow$ IR $\uparrow$ Pick $\uparrow$ CLIP $\uparrow$ FID $\downarrow$
@@ 5 Conclusion
6 Acknowledgment
This work was supported in part by the National Natural Science Foundation of China (Grant 62472098).
@@ Appendix A More Details of Condition Image Token Filtering
We analyze the computational cost of Condition Image Token Filtering (CITF) on three benchmark datasets: COCO-Stuff, ADE20K, and SACap-1M. For each dataset, we consider five settings to comprehensively analyze the computational cost: I) without condition tokens, representing the lowest computation. II–IV) CITF applied with the minimum, maximum, and average numbers of retained condition tokens, respectively. V) without CITF, i.e., no token filtering, representing the highest computation. For each configuration, we report the resulting average image generation time and multiply–accumulate operations (MACs).
@@ Appendix B More Details of Shape Guidance Strength Modulation
Appendix C More Ablation Study Results
C.1 Ablation on AIA
To further verify the effectiveness of the proposed Attribute Isolation Attention Mask (AIA), we conduct an ablation experiment by removing the AIA module from Seg2Any. As illustrated in Figure 8, the absence of AIA leads to attribute leakage across entities.
Figure 8: More qualitative ablation comparisons on the Attribute Isolation Attention Mask (AIA). Without AIA, attribute leakage occurs across entities, e.g., fur texture is transferred between cat breeds (row 1), the red peel of dragon fruit is replaced by kiwi’s green peel (row 2), and the LEGO style spreads to other entities (row 3). In contrast, with AIA, such attribute leakage is effectively eliminated.
C.2 Ablation on SSFA vs. PLACE Attention
To ensure a fair comparison between the proposed Sparse Shape Feature Adaptation (SSFA) and PLACE attention [24], we reimplemented PLACE attention [24] under the same FLUX architecture, while keeping all other training setups fixed. Specifically, we use a learning rate of 1e-4, a batch size of 16, and train for 20k steps on 4 NVIDIA H100 GPUs. As shown in Tables 5 and 6, our model surpasses this PLACE attention variant by ($2.55\%$) MIoU on ADE20K and ($3.36\%$) class-agnostic MIoU on SACap-Eval, confirming the superiority of our design in a strictly fair comparison.
Table 5: Comparison with PLACE attention under the FLUX model on COCO-Stuff and ADE20K.
Method COCO-Stuff ADE20K
MIoU $\uparrow$ FID $\downarrow$ MIoU $\uparrow$ FID $\downarrow$
SAA 43.48 20.57 44.85 33.14
PLACE attention 44.56 20.10 51.56 32.22
SAA+SSFA 45.50 19.06 54.11 32.37
Table 6: Comparison with PLACE attention under the FLUX model on SACap-Eval.
Method
Class-agnostic
MIoU $\uparrow$
Region-wise Quality Global-wise Quality
Spatial $\uparrow$ Color $\uparrow$ Shape $\uparrow$ Texture $\uparrow$ IR $\uparrow$ Pick $\uparrow$ CLIP $\uparrow$ FID $\downarrow$
SAA 89.39 93.36 89.26 85.84 87.86 0.47 21.83 28.09 15.30
PLACE attention 90.8 93.03 89.12 86.36 88.08 0.51 21.90 27.93 15.81
SAA+SSFA 94.16 93.43 89.16 85.61 88.53 0.44 21.81 27.98 15.20
Figure 9 presents the synthesis results on the ADE20K dataset using SAA, PLACE attention, and SAA+SSFA, all trained under the same FLUX architecture. It can be seen that SAA+SSFA generates shapes that best conform to the provided segmentation masks.
Figure 9: More qualitative ablation comparisons on the Sparse Shape Feature Adaptation (SSFA). All methods are trained under the same FLUX architecture.
Appendix D More Qualitative Results
Figure 10: More qualitative results using the Seg2Any method.
We provide additional qualitative results in Figure 10 to illustrate the effectiveness of our approach. Seg2Any shows strong capability in generating high-quality images that faithfully adhere to detailed entity descriptions, enabling flexible and precise control over complex visual attributes such as color, texture, and shape. Figure 11 presents the qualitative results of Seg2Any in densely crowded scenes.
Figure 11: Qualitative results of Seg2Any in densely crowded scenes.
Appendix E More Details on Datasets and Benchmarks
E.1 Comparison with Existing Dense Caption Datasets
Table 7 presents a detailed comparison between our SACap-1M dataset and several recent dense caption datasets. Unlike previous datasets such as PixelLM-MUSE [33], Osprey [51], COCONut-PanCap [8] and Pix2Cap-COCO [50] which are constructed upon closed-set label spaces (e.g., LVIS [12] and COCO [23]), our proposed SACap-1M dataset provides open-set segmented entities, enabling much broader generalization and flexibility in open-set segmentation-mask-to-image generation. Compared to GLaMM-GranD [31], also derived from SA-1B [16], our SACap-1M achieves significantly higher segmentation mask density and caption granularity by employing the advanced open-source Qwen2-VL-72B [43] vision-language model to generate more accurate and fine-grained regional captions.